# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

### v3

@@ 6. Technical Tools: Intersection Joins
Definition 6.1 (Intersection queries (Khamis et al., 2022)).
@@
For the purpose of this paper, we are only interested in a special class of intersection queries, defined below:
@@
Let $\overline{Q}$ be the time extension of a query $Q$. Prior work reduces the evaluation of $\overline{Q}$ to the evaluation of $\widehat{Q}$, where $\widehat{Q}$ is the multivariate extension of $Q$ (Eq. (3)), whose variables are all point variables (Khamis et al., 2022). Moreover, it is shown in (Khamis et al., 2022) that this reduction is optimal (up to a $\widetilde{\mathcal{O}}(1)$ factor). In particular, the query $\overline{Q}$ is exactly as hard as the hardest component $\widehat{Q}_{\bm{\sigma}}$ in $\widehat{Q}$. This is shown by a backward reduction from (the static evaluation of) each component $\widehat{Q}_{\bm{\sigma}}$ to $\overline{Q}$. We explain both reductions in detail in Appendix D and briefly summarize the forward reduction below. We rely on both reductions to prove upper and lower bounds for IVM in the insert-delete setting (Sec. 3.2).
Reduction from $\overline{Q}$ to $\widehat{Q}$
Suppose we want to evaluate $\overline{Q}$ on a given database instance $\overline{\mathcal{D}}$. We construct a segment tree, $\mathcal{T}_{N}$, over the interval variable $[Z]$, where the parameter $N$ is roughly the maximum number of different $Z$-values. A segment tree $\mathcal{T}_{N}$ is a balanced binary tree with $N$ leaves. Each node corresponds to an interval. The root corresponds to the interval $[1,N]$, its left and right child correspond to the intervals $[1,N/2]$ and $[N/2+1,N]$ respectively, and so on. Each node can be represented by a bitstring of length at most $\log_{2}N$. Figure 10 in Appendix D depicts the segment tree $\mathcal{T}_{8}$. Each interval $[z]$ can be broken down into at most $\mathcal{O}(\log N)$ nodes in $\mathcal{T}_{N}$. We refer to these nodes as the canonical partition of $[z]$, and denote them by $\text{\sf CP}_{N}([z])$. For example, the canonical partitions of $[2,8]$ and $[2,5]$ in the segment tree $\mathcal{T}_{8}$ are:
@@
The canonical partition of the intersection of some intervals is roughly the intersection of their canonical partitions (Lemma D.1). Two (or more) nodes in the segment tree correspond to overlapping intervals if and only if the corresponding bitstrings have a prefix-suffix relationship (Lemma D.1). For example, the nodes labeled by $1$ and $100$ in Figure 10 correspond to two overlapping intervals, namely $[5,8]$ and $[5,5]$, because the string $1$ is a prefix of $100$. Therefore, testing intersections can be reduced to testing whether some bitstrings form a chain of prefixes. Armed with this idea, we convert $\overline{\mathcal{D}}$ into a database instance $\widehat{\mathcal{D}}$ for $\widehat{Q}$, called the canonical partition of $\overline{\mathcal{D}}$ and denoted by $\text{\sf CP}_{N}(\overline{\mathcal{D}})$. There is a mapping between the outputs $\overline{Q}(\overline{\mathcal{D}})$ and $\widehat{Q}(\widehat{\mathcal{D}})$ (Theorem D.5).
@@
Suppose we want to reduce $\overline{Q}_{\triangle}$ from Eq. (11) to $\widehat{Q}_{\triangle}$ from Example 2.4. In particular, we are given the database instance $\overline{\mathcal{D}}$ from Figure 2(d) and want to compute the corresponding instance $\widehat{\mathcal{D}}$ for $\widehat{Q}_{\triangle}$. We show how to construct $\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123}$ for $\widehat{Q}_{123}$ in Eq. (4). The remaining $\widehat{Q}_{132},\ldots,\widehat{Q}_{321}$ are similar. The query $\widehat{Q}_{123}$ is meant to test whether $\overline{R},\overline{S},\overline{T}$ contain three intervals $[z^{R}],[z^{S}],[z^{T}]$ whose canonical partitions contain bitstrings $z^{R},z^{S},z^{T}$ where $z^{R}$ is a prefix of $z^{S}$ which is a prefix of $z^{T}$. This can only happen if there are bitstrings $z_{2},z_{3}$ where $z^{S}=z^{R}\circ z_{2}$ and $z^{T}=z^{S}\circ z_{3}$, where $\circ$ denotes string concatenation. We define:
@@
Let $\varepsilon$ denote the empty string. Applying the above to $\overline{\mathcal{D}}$ from Figure 2(d), $\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123}$ respectively contain the tuples: $(\varepsilon,a_{1},b_{1})$, $(\varepsilon,01,b_{1},c_{1})$, $(\varepsilon,01,\varepsilon,a_{1},c_{1})$, which join together producing the output tuple $(\varepsilon,01,\varepsilon,a_{1},b_{1},c_{1})$ of $\widehat{Q}_{123}$. This is a witness to the output tuple $([3,5],a_{1},b_{1},c_{1})$ of $\overline{Q}_{\triangle}$. See Example D.2 for more details.
@@ 7. IVM: Insert-Delete Setting
In this section, we give a brief overview of how we obtain our upper and lower bounds for IVM in the insert-delete setting from Theorems 3.4 and 3.6, respectively. Details are deferred to Appendix E.
@@ 7.1. Upper bound for $\text{IVM}^{\pm}[Q]$
The following example demonstrates our algorithm for $\text{IVM}^{\pm}[Q]$ from Theorem 3.4. Algorithm 2 in Appendix E.1 shows the general algorithm. Appendix E.2 shows how to extend it to $\text{IVM}^{\pm}_{\delta}[Q]$.
Example 7.1 (for Theorem 3.4).
Suppose we want to solve the problem $\text{IVM}^{\pm}[Q_{\triangle}]$ for $Q_{\triangle}$ in Eq. (1), where we have a stream of $N$ single-tuple inserts/deletes into $R,S$ and $T$. For simplicity, assume that $N$ is known in advance (Remark E.1 explains how to remove this assumption.) We introduce a new interval variable $[Z]$ and use it to represent time. Specifically, $[Z]$ will represent the lifespan of every tuple in the database, in the spirit of temporal databases (Jensen and Snodgrass, 2018). By adding $[Z]$ to every atom, we obtain the time extension $\overline{Q}_{\triangle}$ of $Q_{\triangle}$ in Eq. (11). Suppose that the $\tau$-th update is an insert of a tuple $(a,b)$ into $R$, i.e. $+R(a,b)$. Then, we apply the insert $+\overline{R}([\tau,\infty],a,b)$ into the time extension, indicating that the tuple $(a,b)$ lives in $R$ from time $\tau$ on (since we don’t know yet its future deletion time). Now suppose that the $\tau^{\prime}$-th update (for some $\tau^{\prime}>\tau$) is a delete of the same tuple $(a,b)$ in $R$. Then, we replace the tuple $([\tau,\infty],a,b)$ in $\overline{R}$ with $([\tau,\tau^{\prime}],a,b)$. We can extract the result of $Q_{\triangle}$ at any time $\tau$ by selecting output tuples $([z],a,b,c)$ of $\overline{Q}_{\triangle}$ where the interval $[z]$ contains the current time $\tau$. Therefore, if we can efficiently maintain $\overline{Q}_{\triangle}$, we can efficiently maintain $Q_{\triangle}$. Table 2 shows the same stream of 8 updates into $Q_{\triangle}$ as Table 1 but adds the corresponding updates to $\overline{Q}_{\triangle}$. The final database $\overline{\mathcal{D}}$ after all updates have taken place is the same as the one shown in Figure 2(d).
@@ 7.2. Lower bound for $\text{IVM}^{\pm}_{\delta}[Q]$
The following example demonstrates our lower bound for $\text{IVM}^{\pm}_{\delta}[Q]$ from Theorem 3.6, which is based on a reduction from the static evaluation of a component $\widehat{Q}_{\bm{\sigma}}$ of the multivariate extension of $Q$ to $\text{IVM}^{\pm}_{\delta}[Q]$. Algorithm 3 in Appendix E.4 details the reduction.
@@
Consider $Q_{\triangle}$ from Eq. (1) and $\widehat{Q}_{123}$ from Eq. (4). Suppose that there exists $\gamma>0$ such that $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$ can be solved with amortized update time $\widetilde{\mathcal{O}}(N^{\alpha})$ where $N$ is the number of single-tuple updates, $\alpha\stackrel{{\scriptstyle\text{def}}}{{=}}\omega(\widehat{Q}_{123})-1-\gamma$, and $\omega$ is given by Definition 2.2. We show that this implies that $\text{Eval}[\widehat{Q}_{123}]$ can be solved in time $\widetilde{\mathcal{O}}(|\widehat{\mathcal{D}}_{123}|^{\omega(\widehat{Q}_{123})-\gamma}+|\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123})|)$ on any database instance $\widehat{\mathcal{D}}_{123}$, thus contradicting the definition of $\omega(\widehat{Q}_{123})$.
@@
Let us take an arbitrary database instance $\widehat{\mathcal{D}}_{123}=(\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123})$ and evaluate $\widehat{Q}_{123}$ over $\widehat{\mathcal{D}}_{123}$ using the oracle that solves $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$. Example D.6 shows how to reduce $\text{Eval}[\widehat{Q}_{123}]$ over $\widehat{\mathcal{D}}_{123}$ to the static evaluation of $\overline{Q}_{\triangle}$ from Eq. (11) over a database instance $\overline{\mathcal{D}}=(\overline{R},\overline{S},\overline{T})$ that satisfies:
$|\overline{\mathcal{D}}|=|\widehat{\mathcal{D}}_{123}|,\quad\quad|\overline{Q}_{\triangle}(\overline{\mathcal{D}})|=|\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123})|.$
Moreover, $\overline{\mathcal{D}}$ can be constructed in linear time (Theorem D.9.) Next, we reduce the evaluation of $\overline{Q}_{\triangle}$ over $\overline{\mathcal{D}}$ into $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$.
@@
We interpret the interval $[Z]$ of a tuple in $\overline{Q}_{\triangle}$ as the “lifespan” of a corresponding tuple in $Q_{\triangle}$, similar to Section 7.1. In particular, we initialize relations $R,S,T$ for $Q_{\triangle}$. We also sort the tuples $([\alpha,\beta],a,b)$ of $\overline{R}$ based on the end points $\alpha$ and $\beta$ of their $[Z]$-intervals. Each tuple $([\alpha,\beta],a,b)$ appears twice in the order: once paired with its beginning $\alpha$ and another with its end $\beta$. Similarly, we add the tuples $([\alpha,\beta],b,c)$ of $\overline{S}$ and $([\alpha,\beta],a,c)$ of $\overline{T}$ to the same sorted list, where each tuple appears twice. Now, we go through the list in order. If the next tuple $([\alpha,\beta],a,b)$ is paired with its beginning $\alpha$, then we insert the tuple $(a,b)$ into $R$. However, if the next tuple $([\alpha,\beta],a,b)$ is paired with its end $\beta$, then we delete $(a,b)$ from $R$. The same goes for $S$ and $T$. The number of updates is $2|\overline{\mathcal{D}}|$, and the total time needed to process them is $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|^{\alpha+1})$.
@@
To compute the output of $\overline{Q}_{\triangle}$, we do the following. After every insert or delete in $Q_{\triangle}$, we use the oracle for $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$ to enumerate the change in the output of $Q_{\triangle}$. Whenever the oracle reports an insert of an output tuple $(a,b,c)$ at time $\tau$ and the delete of the same tuple at a later time $\tau^{\prime}>\tau$, then we know that $([\tau,\tau^{\prime}],a,b,c)$ is an output tuple of $\overline{Q}_{\triangle}$. Therefore, the overall time needed to compute the output of $\overline{Q}_{\triangle}$ (including the total update time from before) is $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|^{\alpha+1}+|\overline{Q}_{\triangle}(\overline{\mathcal{D}})|)=$ $\widetilde{\mathcal{O}}(|\widehat{\mathcal{D}}_{123}|^{\alpha+1}+|\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123})|)$. This contradicts the definition of $\omega(\widehat{Q}_{123})$.

### v4

@@ 6. Technical Tools: Intersection Joins
Definition 6.1 (Intersection queries (Abo Khamis et al., 2022)).
@@
In this paper, we are only interested in a special class of intersection queries, defined below:
@@
Let $\overline{Q}$ be the time extension of a query $Q$. Prior work reduces the evaluation of $\overline{Q}$ to the evaluation of $\widehat{Q}$, where $\widehat{Q}$ is the multivariate extension of $Q$ (Eq. (7)), whose variables are all point variables (Abo Khamis et al., 2022). Moreover, it is shown in (Abo Khamis et al., 2022) that this reduction is optimal (up to a $\widetilde{\mathcal{O}}(1)$ factor). In particular, the query $\overline{Q}$ is exactly as hard as the hardest component $\widehat{Q}_{\bm{\sigma}}$ in $\widehat{Q}$. This is shown by a backward reduction from (the static evaluation of) each component $\widehat{Q}_{\bm{\sigma}}$ to $\overline{Q}$. We summarize both reductions below and defer the details to Appendix D and (Abo Khamis et al., 2022). We rely on both reductions to prove upper and lower bounds for IVM in the insert-delete setting (Sec. 4.2).
[…]
A segment tree $\mathcal{T}_{8}$. Each node is identified by a bitstring.
6.1. Reduction from $\overline{Q}$ to $\widehat{Q}$
Suppose we want to evaluate $\overline{Q}$ on a given database instance $\overline{\mathcal{D}}$. We construct a segment tree, $\mathcal{T}_{N}$, over the interval variable $[Z]$, where the parameter $N$ is roughly the maximum number of different $Z$-values. A segment tree $\mathcal{T}_{N}$ is a balanced binary tree with $N$ leaves. Each node corresponds to an interval. The root corresponds to the interval $[1,N]$, its left and right child correspond to the intervals $[1,N/2]$ and $[N/2+1,N]$ respectively, and so on. Each node can be represented by a bitstring of length at most $\log_{2}N$. Figure 2 depicts the segment tree $\mathcal{T}_{8}$. Each interval $[z]$ can be broken down into at most $\mathcal{O}(\log N)$ nodes in $\mathcal{T}_{N}$. We refer to these nodes as the canonical partition of $[z]$, and denote them by $\text{\sf CP}_{N}([z])$. For example, the canonical partitions of $[2,8]$ and $[2,5]$ in the segment tree $\mathcal{T}_{8}$ are:
@@ 6.1. Reduction from $\overline{Q}$ to $\widehat{Q}$
The canonical partition of the intersection of some intervals is roughly the intersection of their canonical partitions (Lemma D.1). Two (or more) nodes in the segment tree correspond to overlapping intervals if and only if one node is an ancestor of the other, which can only happen if one of the two corresponding bitstrings is a prefix of the other (Lemma D.1). For example, the nodes labeled by $1$ and $100$ in Figure 2 correspond to two overlapping intervals, namely $[5,8]$ and $[5,5]$, because the string $1$ is a prefix of $100$. Therefore, testing intersections can be reduced to testing whether some bitstrings form a chain of prefixes. Armed with this idea, we convert $\overline{\mathcal{D}}$ into a database instance $\widehat{\mathcal{D}}$ for $\widehat{Q}$, called the canonical partition of $\overline{\mathcal{D}}$ and denoted by $\text{\sf CP}_{N}(\overline{\mathcal{D}})$. There is a mapping between the outputs of $\overline{Q}(\overline{\mathcal{D}})$ and $\widehat{Q}(\widehat{\mathcal{D}})$, as the following example shows. See Theorem D.5 for more details.
@@
Suppose we want to reduce $\overline{Q}_{\triangle}$ from Eq. (14) to $\widehat{Q}_{\triangle}$ from Example 3.4. In particular, we are given the database instance $\overline{\mathcal{D}}$ from Figure 2(d) and want to compute the corresponding instance $\widehat{\mathcal{D}}$ for $\widehat{Q}_{\triangle}$. We show how to construct $\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123}$ for $\widehat{Q}_{123}$ in Eq. (8). The remaining $\widehat{Q}_{132},\ldots,\widehat{Q}_{321}$ are similar. The query $\widehat{Q}_{123}$ is meant to test whether $\overline{R},\overline{S},\overline{T}$ contain three intervals $[z^{R}],[z^{S}],[z^{T}]$ whose canonical partitions contain bitstrings $z^{R},z^{S},z^{T}$ where $z^{R}$ is a prefix of $z^{S}$ which is a prefix of $z^{T}$. This can only happen if there are bitstrings $z_{2},z_{3}$ where $z^{S}=z^{R}\circ z_{2}$ and $z^{T}=z^{S}\circ z_{3}$, where $\circ$ denotes string concatenation. To that end, we define:
@@
Note that $|\widehat{T}_{123}|=\mathcal{O}(|\overline{T}|\cdot\text{\sf polylog}(N))$ and can be constructed in time $\mathcal{O}(|\overline{T}|\cdot\text{\sf polylog}(N))$, because the height of the segment tree is $\mathcal{O}(\log N)$. The same goes for $\widehat{R}_{123}$ and $\widehat{S}_{123}$. Let $\varepsilon$ denote the empty string. Applying the above to $\overline{\mathcal{D}}$ from Figure 2(d), $\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123}$ respectively contain the tuples: $(\varepsilon,a_{1},b_{1})$, $(\varepsilon,01,b_{1},c_{1})$, $(\varepsilon,01,\varepsilon,a_{1},c_{1})$, which join together producing the output tuple $(\varepsilon,01,\varepsilon,a_{1},b_{1},c_{1})$ of $\widehat{Q}_{123}$. This is a witness to the output tuple $([3,5],a_{1},b_{1},c_{1})$ of $\overline{Q}_{\triangle}$. The full answer to $\overline{Q}_{\triangle}$, given by Figure 2(d), can be retrieved from the union of the answers to the six queries $\widehat{Q}_{132},\ldots,\widehat{Q}_{321}$. See Example D.2 for more details.
6.2. Reduction from $\widehat{Q}$ to $\overline{Q}$
We now summarize the backward reduction from the static evaluation of each component $\widehat{Q}_{\bm{\sigma}}$ of $\widehat{Q}$ back to $\overline{Q}$, based on (Abo Khamis et al., 2022). We will utilize this reduction later to establish our lower bound in Sec. 7.2.
Example 6.5.
Continuing with Example 6.3, let us consider the query $\widehat{Q}_{123}$ from (8). Let $\widehat{\mathcal{D}}_{123}^{\prime}$ be an arbitrary database instance for $\widehat{Q}_{123}$. (Unlike Example 6.3, here we cannot make any assumption about how $\widehat{\mathcal{D}}_{123}^{\prime}$ was constructed.) We show how to use an oracle for the intersection query $\overline{Q}_{\triangle}$ from (14) in order to compute $\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123}^{\prime})$ in the same time. WLOG we can assume that each value $x$ that appears in $\widehat{\mathcal{D}}_{123}^{\prime}$ is a bitstring of length exactly $\ell$ for some constant $\ell$. (If a bitstring has length less than $\ell$, we can pad it with zeros.) We construct a segment tree $\mathcal{T}_{N^{\prime}}$ where $N^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}2^{3\ell}$, i.e. a segment tree of depth $3\ell$. (We chose $3$ in this example because it is the number of atoms in $\widehat{Q}_{123}$.) Each node $v$ in the segment tree corresponds to an interval that is contained in $[N^{\prime}]$ and is identified by a bitstring of length at most $3\ell$. Given a bitstring $b$ of length at most $3\ell$, let $\text{\sf seg}_{N^{\prime}}(b)$ denote the corresponding interval in the segment tree $\mathcal{T}_{N^{\prime}}$. We construct the following database instance $\overline{\mathcal{D}}^{\prime}$ for $\overline{Q}_{\triangle}$:
Following (Abo Khamis et al., 2022), we can show that there is a one-to-one mapping between the output tuples of $\overline{Q}_{\triangle}(\overline{\mathcal{D}}^{\prime})$ and the output tuples of $\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123}^{\prime})$. This follows from the observation that the three intervals $\text{\sf seg}_{N^{\prime}}(z_{1}^{R})$, $\text{\sf seg}_{N^{\prime}}(z_{1}^{S}\circ z_{2}^{S})$, $\text{\sf seg}_{N^{\prime}}(z_{1}^{T}\circ z_{2}^{T}\circ z_{3}^{T})$ overlap if and only if $z_{1}^{R}=z_{1}^{S}=z_{1}^{T}$ and $z_{2}^{S}=z_{2}^{T}$. (Recall that the bitstrings $z_{1}^{R},z_{1}^{S},\ldots$ all have the same length $\ell$.) Moreover note that $\overline{R}^{\prime},\overline{S}^{\prime},\overline{T}^{\prime}$ have the same sizes as $\widehat{R}_{123}^{\prime},\widehat{S}_{123}^{\prime},\widehat{\mathcal{T}}_{123}^{\prime}$ respectively and they can be constructed in linear time.
@@ 7. IVM: Insert-Delete Setting
In this section, we give a brief overview of how we obtain our upper and lower bounds for IVM in the insert-delete setting from Theorem 4.4 and 4.5, respectively. Details are deferred to Appendix E.
@@ 7.1. Upper bound for $\text{IVM}^{\pm}[Q]$
In order to prove Theorem 4.4, we start by describing an algorithm that meets a weaker version of the upper bound in Theorem 4.4, where the current database size $|\mathcal{D}^{(\tau)}|$ is replaced by the number of single-tuple updates $N$: (Note that $|\mathcal{D}^{(\tau)}|\leq N$ and could be unboundedly smaller.)
Lemma 7.1.
For any query $Q$, both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(N^{\mathsf{w}(\widehat{Q})-1})$ amortized update time and non-amortized constant enumeration delay, where $\widehat{Q}$ is the multivariate extension of $Q$ and $N$ is the number of single-tuple updates.
In particular, we can show that we can use the algorithm from Lemma 7.1 as a black box in order to meet the stronger upper bound in Theorem 4.4. The main reason why the algorithm from Lemma 7.1 does not immediately meet the upper bound from Theorem 4.4 is that $N$ can grow much larger than the database size $|\mathcal{D}|$, especially in the scenario where the update stream contains many deletes. However, whenever that happens, we can “reset” the algorithm from Lemma 7.1 by restarting from scratch and inserting all the tuples in the current database $\mathcal{D}$ as a stream of $|\mathcal{D}|$ inserts. If this is done carefully, then we can ensure that the total number of updates since the last reset is not significantly larger than $|\mathcal{D}|$. Appendix E.3 gives the details.
We now focus on demonstrating our algorithm that proves Lemma 7.1 for $\text{IVM}^{\pm}[Q]$. Algorithm 2 in Appendix E.1 gives the general algorithm. Appendix E.2 shows how to extend it to $\text{IVM}^{\pm}_{\delta}[Q]$.
Example 7.2 (for Lemma 7.1).
Suppose we want to solve the problem $\text{IVM}^{\pm}[Q_{\triangle}]$ for $Q_{\triangle}$ in Eq. (1), where we have a stream of $N$ single-tuple inserts/deletes into $R,S$ and $T$. For simplicity, assume that $N$ is known in advance. (If $N$ is not known in advance, we can initially assume $N$ to be a constant and keep doubling $N$ every time we exceed it. See Remark E.1 for details.) We introduce a new interval variable $[Z]$ and use it to represent time. Specifically, $[Z]$ will represent the lifespan of every tuple in the database, in the spirit of temporal databases (Jensen and Snodgrass, 2018). By adding $[Z]$ to every atom, we obtain the time extension $\overline{Q}_{\triangle}$ of $Q_{\triangle}$ in Eq. (14). Suppose that the $\tau$-th update is an insert of a tuple $(a,b)$ into $R$, i.e. $+R(a,b)$. Then, we apply the insert $+\overline{R}([\tau,\infty],a,b)$ into the time extension, indicating that the tuple $(a,b)$ lives in $R$ from time $\tau$ on (since we don’t know yet its future deletion time). Now suppose that the $\tau^{\prime}$-th update (for some $\tau^{\prime}>\tau$) is a delete of the same tuple $(a,b)$ in $R$. Then, we replace the tuple $([\tau,\infty],a,b)$ in $\overline{R}$ with $([\tau,\tau^{\prime}],a,b)$. We can extract the result of $Q_{\triangle}$ at any time $\tau$ by selecting output tuples $([z],a,b,c)$ of $\overline{Q}_{\triangle}$ where the interval $[z]$ contains the current time $\tau$. Therefore, if we can efficiently maintain $\overline{Q}_{\triangle}$, we can efficiently maintain $Q_{\triangle}$. Table 2 shows the same stream of 8 updates into $Q_{\triangle}$ as Table 1 but adds the corresponding updates to $\overline{Q}_{\triangle}$. The final database $\overline{\mathcal{D}}$ after all updates have taken place is the same as the one shown in Figure 2(d).
@@ 7.2. Lower bound for $\text{IVM}^{\pm}_{\delta}[Q]$
In order to prove the lower bound from Theorem 4.5, we prove a stronger lower bound where $|\mathcal{D}^{(\tau)}|$ is replaced by the number of single-tuple updates $N$. The following example demonstrates our lower bound for $\text{IVM}^{\pm}_{\delta}[Q]$, which is based on a reduction from the static evaluation of a component $\widehat{Q}_{\bm{\sigma}}$ of the multivariate extension of $Q$ to $\text{IVM}^{\pm}_{\delta}[Q]$. Algorithm 3 in Appendix E.4 details the reduction.
@@
Consider $Q_{\triangle}$ from Eq. (1) and $\widehat{Q}_{123}$ from Eq. (8). Suppose that there exists $\gamma>0$ such that $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$ can be solved with amortized update time $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\kappa})$ where $|\mathcal{D}^{(\tau)}|$ is the database size, $\kappa\stackrel{{\scriptstyle\text{def}}}{{=}}\omega(\widehat{Q}_{123})-1-\gamma$, and $\omega$ is given by Definition 3.2. (Note that $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\kappa})=\widetilde{\mathcal{O}}(N^{\kappa})$ since $|\mathcal{D}^{(\tau)}|\leq N$.) We show that this implies that $\text{Eval}[\widehat{Q}_{123}]$ can be solved in time $\widetilde{\mathcal{O}}(|\widehat{\mathcal{D}}_{123}|^{\omega(\widehat{Q}_{123})-\gamma}+|\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123})|)$ on any database instance $\widehat{\mathcal{D}}_{123}$, thus contradicting the definition of $\omega(\widehat{Q}_{123})$.
@@
Let us take an arbitrary database instance $\widehat{\mathcal{D}}_{123}=(\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123})$ and evaluate $\widehat{Q}_{123}$ over $\widehat{\mathcal{D}}_{123}$ using the oracle that solves $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$. Example 6.5 shows how to reduce $\text{Eval}[\widehat{Q}_{123}]$ over $\widehat{\mathcal{D}}_{123}$ to the static evaluation of $\overline{Q}_{\triangle}$ from Eq. (14) over a database instance $\overline{\mathcal{D}}=(\overline{R},\overline{S},\overline{T})$ that satisfies: $|\overline{\mathcal{D}}|=|\widehat{\mathcal{D}}_{123}|$ and $|\overline{Q}_{\triangle}(\overline{\mathcal{D}})|=|\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123})|$. Moreover, $\overline{\mathcal{D}}$ can be constructed in linear time (Theorem D.8). Next, we reduce the evaluation of $\overline{Q}_{\triangle}$ over $\overline{\mathcal{D}}$ into $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$.
@@
We interpret the interval $[Z]$ of a tuple in $\overline{Q}_{\triangle}$ as the “lifespan” of a corresponding tuple in $Q_{\triangle}$, similar to Section 7.1. In particular, we initialize empty relations $R,S,T$ to be used as input for $Q_{\triangle}$, which will be maintained under a sequence of updates for the $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$ problem. We also sort the tuples $([\alpha,\beta],a,b)$ of $\overline{R}$ based on the end points $\alpha$ and $\beta$ of their $[Z]$-intervals. Each tuple $([\alpha,\beta],a,b)$ appears twice in the order: once paired with its beginning $\alpha$ and another with its end $\beta$. Similarly, we add the tuples $([\alpha,\beta],b,c)$ of $\overline{S}$ and $([\alpha,\beta],a,c)$ of $\overline{T}$ to the same sorted list, where each tuple appears twice. Now, we go through the list in order. If the next tuple $([\alpha,\beta],a,b)$ is paired with its beginning $\alpha$, then we insert the tuple $(a,b)$ into $R$. However, if the next tuple $([\alpha,\beta],a,b)$ is paired with its end $\beta$, then we delete $(a,b)$ from $R$. The same goes for $S$ and $T$. The number of updates, $N$, is $2|\overline{\mathcal{D}}|$, and the total time needed to process them is $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|^{\kappa+1})$.
@@
To compute the output of $\overline{Q}_{\triangle}$, we do the following. After every insert or delete in $Q_{\triangle}$, we use the oracle for $\text{IVM}^{\pm}_{\delta}[Q_{\triangle}]$ to enumerate the change in the output of $Q_{\triangle}$. Whenever the oracle reports an insert of an output tuple $(a,b,c)$ at time $\tau$ and the delete of the same tuple at a later time $\tau^{\prime}>\tau$, then we know that $([\tau,\tau^{\prime}],a,b,c)$ is an output tuple of $\overline{Q}_{\triangle}$. Therefore, the overall time needed to compute the output of $\overline{Q}_{\triangle}$ (including the total update time from before) is $\widetilde{\mathcal{O}}(|\overline{\mathcal{D}}|^{\kappa+1}+|\overline{Q}_{\triangle}(\overline{\mathcal{D}})|)=$ $\widetilde{\mathcal{O}}(|\widehat{\mathcal{D}}_{123}|^{\kappa+1}+|\widehat{Q}_{123}(\widehat{\mathcal{D}}_{123})|)$. This contradicts the definition of $\omega(\widehat{Q}_{123})$.